# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v1

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, yet they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we disclose that CLIP’s vulnerabilities primarily stem from its excessive encoding of class-irrelevant features, which can compromise the model’s visual feature resistivity to input perturbations, making it more susceptible to capturing the trigger patterns inserted by backdoor attacks. Inspired by this finding, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs specially designed deep visual prompt tuning and feature-repelling loss to eliminate excessive class-irrelevant features while simultaneously optimizing cross-entropy loss to maintain clean accuracy. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters relative to CLIP, yet it significantly outperforms state-of-the-art baselines, reducing the attack success rate from 67.53% to 2.76% against SoTA attacks and effectively generalizing its defensive capabilities across multiple datasets. Our anonymous code is available on https://anonymous.4open.science/r/RVPT.
@@ 1 Introduction
Figure 1: The empirical distribution of the cosine similarity between the penultimate-layer features/logits of the original images and those of its slightly, non-adversarially perturbed version. Additionally, we obtain the empirical distribution of the mean cosine similarity between the penultimate-layer features of each image and those of images in its class, referred to as intra-class image-image similarity. Four encoders are evaluated: (a) ViT pre-trained exclusively on ImageNet [40]. (b) ViT from openAI’s CLIP [41]. (c) ViT of the backdoored CLIP tuned with Repulsive Visual Prompt Tuning (RVPT) using 1 shot per class on ImageNet. (d) ViT of the backdoored CLIP tuned with RVPT using 16 shots per class on ImageNet. We also calculate each encoder’s Perturbation Resistivity (PR), which reflects the resistivity of the encoded features against input perturbations, and the Attack Success Rate (ASR) against the backdoored encoder, which indicates the encoder’s vulnerability to backdoor attack during inference. Specifically, CLIP in (c) and (d) is backdoored and attacked by BadCLIP [28].
@@
CLIP’s vulnerabilities to backdoor attacks have prompted the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters [2] or require the availability of poisoned data [57, 19, 56], which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) [21], utilizing only a small portion of downstream clean samples to eliminate the embedded trigger feature directly. However, simply adopting VPT proves ineffective in experiments, as the visual encoder will not extract trigger features from clean samples, making it impossible for VPT to learn to remove them. Notably, we observe that CLIP’s visual features exhibit unusually high sensitivity to even non-adversarial input perturbations, a trait strongly associated with its significant vulnerability to backdoor attacks. Therefore, we can instead leverage VPT to lower this sensitivity and enhance backdoor robustness.
To quantify this sensitivity, we introduce the Perturbation Resistivity (PR) metric. PR is defined as the ratio of the expected cosine similarity between the penultimate-layer features of original and perturbed images and the expected cosine similarity between the logits of these images:
$\text{PR}(D,\boldsymbol{\delta})=\frac{\mathbb{E}_{(\boldsymbol{x},y)\sim D}[\cos(\phi(\boldsymbol{x}),\phi(\boldsymbol{x}+\boldsymbol{\delta}))]}{\mathbb{E}_{(\boldsymbol{x},y)\sim D}[\cos(g(\boldsymbol{x}),g(\boldsymbol{x}+\boldsymbol{\delta}))]}.$ (1)
Here, $\phi(\boldsymbol{x})$ represents the penultimate-layer features for the input $\boldsymbol{x}$, $g(\boldsymbol{x})$ denotes the logits for the input $\boldsymbol{x}$, $D$ represents the given dataset and $\cos(\cdot,\cdot)$ denotes cosine similarity. A lower PR value indicates a greater sensitivity of the encoded features to input perturbations, reflecting the model’s potential vulnerability against backdoor attacks due to increased susceptibility to capturing trigger patterns. In particular, we measured the ASRs for models with different PRs and observed a strong negative correlation between ASR and PR, which is presented in the supplemental material. Notably, unless otherwise noted, we abbreviate empirical PR as PR.
In order to investigate the PR values of different visual encoders, we randomly sampled 5,000 images from the ImageNet [9] validation set and introduced random noise [15] of 16$\times$16 pixels as the perturbation to these images. Results for more perturbations are provided in the supplemental material. We expect all the PR values to be close to 1 since the perturbation is neither adversarial nor semantic. However, the results depicted in Figure 1 (a) (b) show that the PR value of CLIP ViT (Vision Transformer [11]) is greatly lower than that of ViT trained on ImageNet, indicating CLIP’s high sensitivity to perturbations and vulnerability to backdoor attacks. To explain this phenomenon, we also analyze each model’s intra-class image-image similarity of the sampled images in Figure 1 and discover that CLIP demonstrates lower similarity compared with traditional supervised models, suggesting that CLIP encodes a broader range of class-irrelevant features, which do not correlate with any class in the specific downstream dataset. Therefore, CLIP’s tendency to excessively encode class-irrelevant features is supposed to be a key factor contributing to its lower PR, as it enables CLIP to capture and encode diverse and fine-grained patterns more easily. In summary, the unique training paradigm of CLIP results in its capture of a broad and potentially excessive range of class-irrelevant features, thereby heightening the sensitivity of its visual features to input perturbations and rendering the CLIP model vulnerable to backdoor attacks.
@@
To address this challenge, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), to trim the class-irrelevant features (CIFs) in a few-shot manner. As shown in Figure 2, RVPT mainly utilizes deep visual prompt tuning to eliminate CIFs. Specifically, it adopts feature-repelling loss to repel the prompted features from the original features by minimizing their mean cosine similarity while optimizing cross-entropy loss to maintain clean accuracy. This mechanism ensures that CIFs, which do not contribute to the cross-entropy loss, are minimized, thereby enhancing the model’s robustness against potential backdoor attacks. As a result, Figure 1 (c) and (d) shows that RVPT effectively increases PR by reducing class-irrelevant features, hence successfully defending against state-of-the-art multimodal backdoor attack, even with one shot per class. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 2 Related Work
Backdoor attacks and defenses on supervised learning. Backdoor attacks have become an increasingly serious security issue because more and more practitioners choose to adopt third-party datasets, platforms, or even backbones to reduce costs. Research on backdoor attacks has primarily focused on designing triggers that enhance stealthiness [6, 46] and effectiveness [31, 36]. For backdoor defenses, researchers employ various strategies to mitigate backdoor attacks, including: (1) pre-processing defense [45] (2) pre-training defense [5] (3) post-training defense [63, 51] (4) test-time defense [14]. Specifically, our work aligns with the post-training defense category, which assumes that defenders do not have access to the poisoned training data and must eliminate the backdoor threat that has already been implanted into the model using only a limited number of clean samples. There are two main strategies for post-training defense. (1) Pruning-based Methods [30, 52]: These approaches focus on identifying the most suspicious neurons and removing them. (2) Fine-tuning-based Methods [49, 63, 35, 26]: These methods purify the compromised model by directly tuning it with clean data through various algorithms.
@@
Backdoor attacks and defenses on CLIP. Although vision-language models (VLMs) [41, 20, 29] have demonstrated significant improvements across various fields, recent research [3, 58] has revealed vulnerabilities of contrastively pre-trained VLMs to backdoor attacks, drawing substantial attention in the field of backdoor attacks and defenses on CLIP-like models. In terms of attacks, BadCLIP [28] optimizes visual trigger patterns within a dual-embedding guided framework, rendering the attack undetectable. Moreover, many backdoor attack methods on multimodal large language models [32, 33, 27, 38, 34] have been proposed recently. Defense strategies can be primarily categorized into three temporal categories: pre-training defense and post-training defense. For pre-training defense [57, 19, 56], SAFECLIP [57] prevents learning from poisoned examples by dynamically dividing the training data into the safe and risky sets by cosine similarity and only applying uni-modal contrastive loss to different modalities of the risky set. For fine-tuning defense [2, 55, 22], CleanCLIP [2] reduces the impact of spurious relationships of backdoor attacks by enforcing the model to realign representations for independent modalities.
@@
Prompt learning for vision-language models.  Lately, various fine-tuning methods have been introduced to further enhance the impressive capabilities of vision-language models (VLMs) across diverse downstream tasks [23, 18, 62, 54], with prompt learning [61, 42, 59] emerging as a particularly efficient and effective approach. Specifically, in the area of natural language processing, prompt learning regards the text prompt, e.g., “a photo of a <CLS>.” as continuous learnable parameters and optimizes it using only a few examples from the downstream dataset. CoOp [61] first introduces prompt learning to VLMs and proposes a method that optimizes the context of the text prompt while keeping the class name tokens fixed. VPT [21] adjusts the visual features by concatenating the image or image features with the learnable prompt. Besides, BadCLIP [1] injects backdoor during the prompt learning stage, enabling the malicious alteration of text features through trigger-aware prompts for a powerful attack. Despite the popularity and versatility of prompt learning for enhancing the model’s robustness [22, 24, 7, 60, 10], limited research has focused on its application in defending against multimodal backdoor attacks. In particular, current post-training backdoor defense methods for VLM mainly focus on tuning model parameters to mitigate spurious relationships, neglecting the potential of prompt learning. Therefore, We propose Repulsive Visual Prompt Tuning, aiming to fill the research gap by incorporating learnable parameters into the features to defend against multimodal backdoor attacks while freezing the model’s parameters.
@@ 3 The Proposed Approach
Figure 2: Illustration of RVPT. RVPT concatenates features of later layers with tunable visual prompts while keeping the CLIP’s pre-trained parameters frozen. To optimize the visual prompts, RVPT employs both cross-entropy (CE) loss and feature-repelling (FR) loss. The FR loss minimizes the mean cosine similarity between the prompted features and the original features, effectively filtering out class-irrelevant features that do not contribute to the CE loss. Meanwhile, the CE loss ensures the clean accuracy of the prompted model.
@@ 3.1 Basic Settings
Threat model.  We assume the adversary has access to the training dataset and can produce a backdoored model by injecting carefully crafted poisoned data into the original clean dataset. For multimodal attacks, the poisoned data usually consists of images with the trigger pattern and the proxy captions for the target class. Once trained on the poisoned dataset, the model misclassifies inputs containing the specific trigger pattern into the predefined target class while maintaining correct classification for all other benign inputs. Afterward, the adversary may release the backdoored model online, raising significant security concerns.
@@ 3.2 Backdoor Attacks on CLIP
In general, the CLIP model comprises a visual encoder $\mathcal{V}(\cdot)$, a textual encoder $\mathcal{T}(\cdot)$, and a projection matrix $\boldsymbol{P}$ that projects both the visual and textual representations into the joint feature space. The training dataset contains image-text pairs represented as $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{t}_{i}\right)\right\}_{i=1}^{N}$, where $\boldsymbol{t_{i}}$ serves as a short caption for the corresponding image $\boldsymbol{x_{i}}$.
During the attack on CLIP in the training stage, adversaries typically inject a small portion of poisoned image-text pairs denoted as $\widetilde{\mathcal{D}_{p}}=\{(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{i})\}_{i=1}^{M}$, with $\widetilde{\boldsymbol{x}}_{i}=\boldsymbol{x_{i}}+\Theta$. $\widetilde{\boldsymbol{x}}_{i}$ represents a poisoned image obtained by modifying the original image $\boldsymbol{x}_{i}$ with the trigger pattern $\Theta$, while $\widetilde{t}_{i}$ denotes the proxy caption for the target class $y_{t}$. Consequently, the original benign training dataset $\mathcal{D}$ can be poisoned, resulting in $\widetilde{\mathcal{D}}=\{\widetilde{\mathcal{D}_{p}}\cup\mathcal{D}\}$. During training, CLIP employs contrastive learning to maximize the cosine similarity of positive pairs, expressed as $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{i})=\cos(\boldsymbol{P}\mathcal{V}(\widetilde{\boldsymbol{x}}_{i}),\boldsymbol{P}\mathcal{T}(\widetilde{\boldsymbol{t}}_{i}))$, and minimize the similarity of negative pairs $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{j\neq i})$. To clarify, the selected training image-text pairs $(\widetilde{\boldsymbol{x}},\widetilde{\boldsymbol{t}})$ belong to $\widetilde{\mathcal{D}}$.
@@ 3.3 Repulsive Visual Prompt Tuning
In this part, we present Repulsive Visual Prompt Tuning (RVPT), a streamlined yet powerful framework designed to defend against multimodal backdoor attacks. Our key insight is to reduce trigger features of the attacked images by restraining the encoding of class-irrelevant features at each layer of the CLIP visual encoder. This approach will increase CLIP’s perturbation resistivity, which is positively correlated with CLIP’s backdoor robustness.
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP, whose visual encoder is Vision Transformer [11]. For the clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is defined as $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings.
Assume the visual encoder has $L$ layers, then starting from the $D$-th layer, we introduce $L-D$ repulsive visual prompts, each of length $b$, defined as $\{\boldsymbol{p}^{k}\in\mathbb{R}^{b\times d_{v}}\}_{d=D}^{L-1}$. These prompts are concatenated with the features from each subsequent block recursively, following the formulation:
@@
Additionally, we also optimize the repulsive prompt with the cross-entropy (CE) loss to avoid repelling the class-relevant features and maintain clean accuracy, formulated as:
@@
where $\tau$ denotes a temperature parameter to control the logit distribution, $T(\cdot)$ is the proxy caption of the input class, e.g., “a photo of a <CLS>.” and $\widetilde{\phi}(\boldsymbol{x}_{i},\boldsymbol{t}_{j})=\cos(\boldsymbol{P}\widetilde{\mathcal{V}}(\boldsymbol{x}_{i}),\boldsymbol{P}\widetilde{\mathcal{T}}(\boldsymbol{t}_{j}))$ is the function of cosine similarity computed with outputs by the recovering encoders.
@@
For further explanation, RVPT can be interpreted through the lens of natural selection [8]. To be specific, FR loss creates a challenging environment for visual prompt tuning to optimize CE loss. In order to simultaneously minimize CE loss and maximize the discrepancy between the prompted features and original features, the visual prompts must learn to select the most suitable class-relevant features while discarding those that are class-irrelevant. Ultimately, this process leads to a more compact visual encoding mechanism.

### v2

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, while they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we reveal that CLIP’s vulnerabilities primarily stem from its tendency to encode features beyond in-dataset predictive patterns, compromising its visual feature resistivity to input perturbations. This makes its encoded features highly susceptible to being reshaped by backdoor triggers. To address this challenge, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs deep visual prompt tuning with a specially designed feature-repelling loss. Specifically, RVPT adversarially repels the encoded features from deeper layers while optimizing the standard cross-entropy loss, ensuring that only predictive features in downstream tasks are encoded, thereby enhancing CLIP’s visual feature resistivity against input perturbations and mitigating its susceptibility to backdoor attacks. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters in CLIP, yet it significantly outperforms state-of-the-art defense methods, reducing the attack success rate from 89.70% to 2.76% against the most advanced multimodal attacks on ImageNet and effectively generalizes its defensive capabilities across multiple datasets. Our code is available on https://github.com/zhangzf01/RVPT.
@@ 1 Introduction
Figure 1: Perturbation resistivity (PR) across different layers of the encoders under various perturbations, including the trigger pattern. A higher PR value indicates less sensitivity to input perturbation. Four encoder variants are evaluated: (a) ViT trained exclusively on clean ImageNet [44]. (b) ViT in backdoored CLIP [45]. (c) ViT in backdoored CLIP tuned with Visual Prompt Tuning (VPT) on ImageNet. (d) ViT of the backdoored CLIP tuned with our proposed RVPT on ImageNet. Specifically, The trigger of the backdoored CLIP is BadCLIP [31]. Detailed experimental settings and additional PR results for CLIP backdoored by other triggers can be found in the supplementary material.
@@
CLIP’s vulnerabilities to backdoor attacks have inspired the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters [3] or require the availability of poisoned data [61, 22, 60], which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) [24], utilizing only a small portion of downstream clean samples to encourage the model to ignore the trigger feature in the poisoned images. However, simply adopting VPT proves ineffective experimentally, as the trigger feature cannot be extracted from clean samples, preventing VPT from learning how to disregard them. Alternatively, VPT can help CLIP learn to extract in-dataset predictive features from clean samples using cross-entropy loss. If the visual encoder is guided to focus exclusively on these features, CLIP will gain backdoor robustness since the trigger features learned in pre-training are ignored during backdoor attacks.
Nevertheless, we observe that CLIP’s encoded visual features, even after VPT, remain highly sensitive to perturbations, which suggests that it still encodes features beyond in-dataset predictive features, leaving it susceptible to potential backdoor attacks. To quantify this sensitivity, we introduce the Perturbation Resistivity (PR) metric, defined as the cosine similarity between the encoded feature $f^{v}(\boldsymbol{x})$ of the image $\boldsymbol{x}$ and the encoded feature $f^{v}(\boldsymbol{x}+\boldsymbol{\delta})$ of its perturbed version: $\boldsymbol{x}+\boldsymbol{\delta}$, where $f^{v}(\cdot)$ is the visual feature extractor:
$\text{PR}(\boldsymbol{x},\boldsymbol{\delta})=\cos(f^{v}(\boldsymbol{x}),f^{v}(\boldsymbol{x}+\boldsymbol{\delta})).$ (1)
A lower PR value indicates greater sensitivity of the encoded features to input perturbations, making the model more susceptible to backdoor triggers. To illustrate this vulnerability of CLIP, we randomly sample 1,000 images from ImageNet and apply various perturbations. We then compare the average PR values of these images across different model variants. As shown in Figure 1 (b) (c), the deeper visual layers of CLIP exhibit significantly higher sensitivity to input perturbations, even after VPT, which allows the trigger to maliciously reshape the encoded visual features layer by layer, enabling successful attacks. In contrast, Figure 1 (a) shows that the ViT supervised on clean ImageNet exhibits lower sensitivity, because it learns only from ImageNet, preventing it from encoding irrelevant features beyond its dataset. Therefore, to reduce CLIP’s sensitivity to input perturbations and enhance its backdoor robustness, we should suppress irrelevant features learned during pre-training and guide it to focus on encoding in-dataset predictive features.
@@
To address this challenge, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), which improves CLIP’s PR by allowing it to exclusively encode the predictive features in the downstream tasks in a few-shot manner. As illustrated in Figure 2, RVPT jointly optimizes two objectives: (1) a multi-layer feature-repelling loss that minimizes the cosine similarity between prompted features and original CLIP features at each layer, and (2) cross-entropy loss for downstream task accuracy. This mechanism forces the model to only extract features that contribute to CE loss optimization during training, while ignoring non-essential features (e.g., noise or trigger pattern). As a result, Figure 1 (d) shows that RVPT significantly improves the PR of CLIP on ImageNet, hence effectively defending against the state-of-the-art multimodal backdoor attack. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 2 Related Work
Backdoor attacks and defenses on supervised learning. Backdoor attacks have become an increasingly serious security issue because more and more practitioners choose to adopt third-party datasets, platforms, or even backbones to reduce costs. Research on backdoor attacks has primarily focused on designing triggers that enhance stealthiness [7, 50] and effectiveness [34, 40]. For backdoor defenses, researchers employ various strategies to mitigate backdoor attacks, including: (1) pre-processing defense [49] (2) pre-training defense [6] (3) post-training defense [68, 55] (4) test-time defense [15]. Specifically, our work aligns with the post-training defense category, which assumes that defenders do not have access to the poisoned training data and must eliminate the backdoor threat that has already been implanted into the model. There are two main strategies for post-training defense. (1) Pruning-based Methods [33, 56]: These approaches focus on identifying and removing the most suspicious neurons. (2) Fine-tuning-based Methods [53, 68, 39, 29]: These methods purify the backdoored model by directly tuning its parameters with clean data.
@@
Backdoor attacks and defenses on CLIP. Although vision-language models (VLMs) [45, 23, 32] have demonstrated significant improvements across various fields, recent research [4, 62] has revealed vulnerabilities of contrastively pre-trained VLMs to backdoor attacks, drawing substantial attention in the field of backdoor attacks and defenses on CLIP-like models. In terms of attacks, BadCLIP [31] optimizes visual trigger patterns within a dual-embedding guided framework, rendering the attack effective and undetectable. Moreover, many backdoor attack methods on multimodal large language models [35, 36, 30, 42, 37] have been proposed recently. Defense strategies can be primarily categorized into two categories: pre-training defense and post-training defense. For pre-training defense [61, 22, 60], SAFECLIP [61] prevents learning from poisoned examples by dynamically dividing the training data into the safe and risky sets by cosine similarity and only applying uni-modal contrastive loss to different modalities of the risky set. For fine-tuning defense [3, 59, 25], CleanCLIP [3] reduces the impact of spurious relationships of backdoor attacks by enforcing the model to realign representations for independent modalities.
@@
Prompt learning for vision-language models.  Lately, various fine-tuning methods have been introduced to further enhance the impressive capabilities of vision-language models (VLMs) across diverse downstream tasks [26, 20, 67, 58], with prompt learning [66, 46, 64] emerging as a particularly efficient and effective approach. Specifically, CoOp [66] first introduces prompt learning to VLMs and proposes a method that optimizes the context of the text prompt, e.g., “a photo of a <CLS>.” while keeping the class name tokens fixed. VPT [24] adjusts the visual features by concatenating the image or image features with the learnable prompt. Besides, BadCLIP [1] injects backdoor during the prompt learning stage, enabling the malicious alteration of text features through trigger-aware prompts for a powerful attack. We notice that despite the effectiveness of prompt learning for enhancing the model’s robustness [25, 27, 8, 65, 11], limited research has explored its use in defending against multimodal backdoor attacks. In particular, current post-training backdoor defense methods for VLMs mainly focus on fine-tuning the whole backdoored models, overlooking the potential of prompt learning. To address the gap, we propose Repulsive Visual Prompt Tuning, which incorporates learnable parameters into features to defend against multimodal backdoor attacks while freezing the model’s parameters.
@@ 3 The Proposed Approach
Figure 2: Illustration of RVPT. RVPT concatenates the features from later layers with tunable visual prompts while keeping the original parameters of CLIP frozen. To optimize the visual prompts, RVPT employs both cross-entropy (CE) loss and feature-repelling (FR) loss. The FR loss minimizes the mean cosine similarity between the prompted features and the original features across layers. Meanwhile, the CE loss ensures the clean accuracy of the prompted model. Together, these losses guide the model to encode only in-dataset predictive features that contribute to CE loss optimization, thereby enhancing its backdoor robustness.
@@ 3.1 Basic Settings
Threat model.  We assume the adversary has access to the training dataset and can produce a backdoored model by injecting carefully crafted poisoned data into the original clean dataset. For multimodal attacks, the poisoned data usually consists of images with the trigger pattern and the proxy captions for the target class. Once trained on the poisoned dataset, the model will misclassify inputs containing the trigger pattern into the predefined target class while maintaining correct classification for all other benign inputs.
@@ 3.2 Backdoor Attacks on CLIP
In general, the CLIP model comprises a visual encoder $\mathcal{V}(\cdot)$, a textual encoder $\mathcal{T}(\cdot)$, and projection matrices that project the encoded representations into a joint feature space : $\boldsymbol{P}_{\mathrm{I}}$ for the image modality and $\boldsymbol{P}_{\mathrm{T}}$ for the text modality. The training dataset contains $N_{1}$ image-text pairs represented as $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{t}_{i}\right)\right\}_{i=1}^{N_{1}}$, where $\boldsymbol{t}_{i}$ serves as a short caption for the corresponding image $\boldsymbol{x}_{i}$.
During the backdoor attack on CLIP in the training stage, adversaries typically inject a small number of $N_{2}$ poisoned image-text pairs denoted as $\mathcal{D}_{p}=\{(\boldsymbol{x}^{p}_{i},\boldsymbol{t}^{p}_{i})\}_{i=1}^{N_{2}}$. $\boldsymbol{x}^{p}_{i}$ represents a poisoned image obtained by modifying the original image $\boldsymbol{x}_{i}$ with the trigger pattern $\Theta$, while $t^{p}_{i}$ denotes the proxy caption for the target class $y_{t}$. Consequently, the original benign training dataset $\mathcal{D}$ can be poisoned, resulting in: $\widetilde{\mathcal{D}}=\{\mathcal{D}\cup\mathcal{D}_{p}\}$. During training, CLIP is optimized on $\widetilde{\mathcal{D}}$ to maximize the cosine similarity of positive pairs, expressed as $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{i})=\cos(\boldsymbol{P}_{\mathrm{I}}\mathcal{V}(\widetilde{\boldsymbol{x}}_{i}),\boldsymbol{P}_{\mathrm{T}}\mathcal{T}(\widetilde{\boldsymbol{t}}_{i}))$, and minimize the similarity of negative pairs $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{j\neq i})$.
@@ 3.3 Repulsive Visual Prompt Tuning
In this part, we present Repulsive Visual Prompt Tuning (RVPT), a framework designed to defend against multimodal backdoor attacks. Our key insight is to make CLIP ignore trigger features of the attacked images by restraining CLIP’s visual encoder to only encode in-dataset predictive features.
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP with the Vision Transformer [12] as its visual encoder. For clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will finally be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings. Specifically, $M$ and $d_{v}$ are hyperparameters representing the number of patches and dimensions of the visual embedding.
Assume the visual encoder has $L$ layers, then starting from the $D$-th layer, we introduce $L-D$ repulsive visual prompts, each of length $b$, defined as $\{\boldsymbol{p}^{d}\in\mathbb{R}^{b\times d_{v}}\}_{d=D}^{L-1}$. These prompts are concatenated with the features from each subsequent block recursively, following the formulation:
@@
Additionally, we also optimize the prompts with the cross-entropy (CE) loss to avoid repelling the in-dataset predictive features and maintain clean accuracy, formulated as:
@@
where $\tau$ denotes a temperature parameter to control the logit distribution and is set as specified in [45], $T(\cdot)$ is the proxy caption of the input class, e.g., “a photo of a <CLS>.” and $\widetilde{\phi}(\boldsymbol{x}_{i},\boldsymbol{t}_{j})=\cos(\boldsymbol{P}_{\mathrm{I}}\widetilde{\mathcal{V}}(\boldsymbol{x}_{i}),\boldsymbol{P}_{\mathrm{T}}\widetilde{\mathcal{T}}(\boldsymbol{t}_{j}))$ means cosine similarity of the outputs from the recovering encoders.
@@
For further explanation, RVPT can be interpreted through the lens of natural selection [9]. To be specific, FR loss creates a challenging environment for visual prompt tuning to optimize CE loss. In order to simultaneously minimize CE loss and maximize the discrepancy between the prompted features and original features, the visual prompts must learn to select the most suitable features to optimize CE loss while ignoring those that are non-essential. Ultimately, this process leads to a more compact visual encoding mechanism, enhancing CLIP’s perturbation resistivity to trigger patterns.